# One Battle After Another: Probing LLMs’ Limits on Multi-Turn Instruction Following with a Benchmark Evolving Framework

### v2

@@ Abstract
Evaluating LLMs’ instruction-following ability in multi-topic dialogues is essential yet challenging. Existing benchmarks are limited to a fixed number of turns, susceptible to saturation and failing to account for users’ interactive experience. In this work, we propose a novel framework backed by a three-layer tracking mechanism and a query synthesis agent to mimic sequential user behaviors. Incorporating Flow Theory, we introduce process-centric metrics and terminate a conversational evaluation only upon exhausting user patience. Upon this framework, we present EvolIF, an evolving benchmark covering 12 constraint groups. Results indicate that GPT-5 excels, sustaining 14 turns with 66.40% robustness. It outperforms Gemini-3.0-Pro by a margin of 5.59%, while other models trail behind. Resources are available at https://github.com/JiaQiSJTU/EvolvingInstructionFollowing.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has catalyzed the development of increasingly sophisticated applications, ranging from extended conversational systems rakotonirina-etal-2025-tools to autonomous agent frameworks hu2025evaluating. The efficacy of these systems is fundamentally predicated on an LLM’s ability to consistently adhere to instructions throughout conversations spanning multiple topics with evolving constraints. This core capability demands robust long-context processing and stateful memory management. Consequently, designing evaluation frameworks for multi-turn instruction following has emerged as a critical research focus he2024multi; kwan2024mt; li-etal-2025-structflowbench.
@@
IFEval zhou2023instruction 1
ComplexBench wen2024benchmarking 1
Multi-IF he2024multi 3
@@
Meeseeks wang2025ask 3
EIFBENCH zou2025eifbenchextremelycomplexinstruction 1
@@
Existing benchmarks suffer from limitations that impede effective evaluation, as exemplified in Fig. 1. First, they fail to capture the interaction dynamics hao2024meta; zhang2025delrec and extended duration typical of real-wolrd scenarios. As shown in Table 1, most benchmarks are restricted to a short interaction window, predominately fewer than 7 turns kwan2024mt, and neglect scenarios involving interleaved topics he2024multi; fan2025fairmtbench. Second, their static nature leads to rapid performance saturation. As LLMs advance, fixed benchmark challenges are quickly mastered he2024multi; bai-etal-2024-mt. Although some benchmarks offer adjustable complexity li2025mtr, maintaining challenge levels via continuous sample generation incurs prohibitive computational costs for model re-evaluation. Third, current methodologies overlook the process-centric aspects of user experience. Inheriting the paradigm from single-turn tasks zhou2023instruction; zhang-etal-2025-cfbench, these benchmarks prioritize final-answer accuracy he2024multi; li-etal-2025-structflowbench; wang2025ask. They neglect interaction stability and fail to provide a direct indication of the maximum number of turns that LLMs can maintain high-fidelity instruction following.
To overcome these shortcomings, we propose a novel and extensible framework for the dynamic generation and process-centric evaluation of complex multi-turn dialogues. Our approach decouples user queries into underlying intentions and surface form. Intention is tracked via a three-layer mechanism that simulates dynamic user behaviors, while the surface form is synthesized by an agent equipped with an LLM-based generator and rigorous validity checkers. Besides, we move beyond single-turn accuracy to emphasize process-centric experience, drawing upon the Flow Theory csikszentmihalyi1990flow. We introduce the notion of patience to model user stickiness to a conversation, where consecutive frustrations lead to a dialogue termination. And we define a suite of process-centric metrics to quantify user experience such as endurance and robustness.
Leveraging this framework, we introduce EvolIF, a benchmark grounded on 541 topics, 12 groups of commonly-adopted constraint groups and 500 diverse user styles. Through an evaluation of 10 leading LLMs, we observe a distinct performance stratification. GPT-5 and Gemini-2.5-Pro establish a commanding lead, whose process-centric scores are nearly double or triple of open-source models. Besides, LLMs share a common and steepest performance drop at around turn 4 and 11, revealing critical bottlenecks in their ability to manage accumulated constraints and complex state transitions.
@@ 2.1 Multi-turn Dialogue Benchmarks
First, script-based evaluations li-etal-2025-structflowbench; deshpande-etal-2025-multichallenge; jia2025simulbench utilize static conversational logs, derived either from human-bot interactions or simulated histories, to assess a model’s response to the final user query. While this approach ensures controlled and consistent LLM comparison, it fails to capture the interactive nature of dialogue, where a model’s prior responses fundamentally influence the conversational trajectory.
Second, a line of work employs pre-defined templates zheng2023judging; fan2025fairmtbench; han2025can. This approach is labor-intensive, requiring significant human effort to design fixed user query sequences. Consequently, these benchmarks face scalability limitations regarding conversational depth and are susceptible to saturation as models become overly optimized to the test set over time.
Third, researchers have explored using LLMs as user simulators zhu2024reliable; sekulic-etal-2024-reliable and evaluation methods based on conversations between LLMs duan-etal-2024-botchat; zhao-etal-2025-auto. Nevertheless, such interactions are prone to uncontrolled divergence and exhibit inherent biases, such as family bias wataoka2025selfpreferencebiasllmasajudge.
@@ 2.2 Instruction Following Benchmarks
One line of work assesses models’ capabilities within increasingly intricate single-turn interaction. Early benchmarks like CIF li-etal-2024-cif evaluated a single constraint per instruction. Subsequent work has evolved to incorporate multiple constraints zhou2023instruction; jiang-etal-2024-followbench; wen2024benchmarking; he2024can or multiple tasks chen-etal-2024-sifo; zou2025eifbenchextremelycomplexinstruction.
A parallel stream of work benchmarks models’ instruction adherence across multiple turns. Multi-IF he2024multi extends IFEval to 3 turns, while MultiTurnInstruct han2025can employs pre-defined templates for diverse scenarios. StructFlowBench li-etal-2025-structflowbench leverages 6 structure types to curate complex dialogue histories. Other studies focus on specialized abilities, such as self-correction wang2025ask, or domain-specific tasks like code generation wang2025codeif.
@@ 3.1 Overview
Flow theory csikszentmihalyi1990flow points out that users enter a psychological state of immersion which can withstand minor disruptions but collapses under prolonged failure. Underpinned by this theory, we employ an adaptive evaluation protocol where the length of a dialogue is contingent on model performance, governed by a “patience” threshold. High-performing models face progressively longer and more challenging threads, while repeated failures will deplete patience and trigger termination. In this way, our benchmark remains a persistent challenge for advanced models, resisting from saturation.
@@ 3.3 Evaluation Protocol
Our evaluation protocol is adaptive and designed to mirror real-world user interactions, premised on Flow Theory csikszentmihalyi1990flow and the cooperative principles of dialogue grice1975logic. Repeated failures by a conversation partner serve as a primary catalyst for user frustration, leading to the eventual disengagement ang2002prosody; hernandez-caralt-etal-2025-stupid.
@@
$P_{t}=\begin{cases}P_{t-1}-1,&\text{if succeed,}\\ P_{max},&\text{otherwise.}\end{cases}$ (3)
@@ 3.4 Evaluation Metrics
Conventional metrics, such as Constraint Satisfaction Rate (CSR) and Instruction Satisfaction Rate (ISR)zhang-etal-2025-cfbench; li-etal-2025-structflowbench, focus primarily on outcome accuracy. To capture the nuances of the conversational process, we introduce a suite of process-centric metrics. Given a benchmark of $D$ dialogues, these metrics are defined below (see AppendixA for details).
@@ 4.1 EvolIF Benchmark
We first curated its core assets: topics, constraints and styles. We collected 541 dialogue topics from IFEval zhou2023instruction, manually removing the attached constraints to isolate the core task scenarios and subjects. To support our dynamic generation process, we assigned a set of customized keywords for each topic. Concurrently, we consolidated constraints from prior works zhou2023instruction; li-etal-2025-structflowbench and our own construction, and systematically re-categorized them into 12 mutually exclusive groups based on semantic intention. These contain 9 objective constraints assessed by rules and 3 subjective constraints measured with an LLM judge. Moreover, we gathered 500 styles by prompting GPT-4.1 with personas from nvidia2025. More in Appendix B.
We guarantee the quality of the benchmark through the following considerations. To ensure integrity and complexity, we applied an automated filter to discard trivial samples, removing dialogues where the average number of constraints over the first 20 turns was less than two. To mitigate family bias spiliopoulou2025play introduced by a single synthesizer, we adopted GPT-4.1, Gemini-2.5-Flash and DeepSeek-V3.1 as synthesizers to generate dialogue sessions with $k=3$ trials.
@@ 4.2 Evaluated Models
We conducted evaluation on ten state-of-the-art large language models from different institutions. They include GPT-5-2025-08-07, Gemini-3.0-Pro-Exp comanici2025gemini, DeepSeek-V3.2-Exp liu2024deepseek, Kimi-K2-Instruct-0905 team2025kimi, Qwen-235B-A22B-Instruct-2507 yang2025qwen3, Grok-4-Fast-Reasoning, Llama-4-Maverick, Seed-1.6-Thinking-250715, MiniMax-M2 and Mistral-Large-3. All of the models were evaluated through API calls with corresponding default settings.
@@ 5.1 Main Results
The performance of LLMs on EvolIF is presented in Table 2. Our analysis reveals a distinct stratification in multi-turn instruction-following capabilities. GPT-5 establishes itself as the state-of-the-art, with Gemini-2.5-Pro following closely behind. These two models demonstrate a superior performance, achieving process-centric scores that are double or even triple of subsequent models. MiniMax-M2 emerges as the most competitive open-source LLMs, forming a second tier alongside Kimi-K2, Qwen3-235B and Grok-4-Fast. The rest constitute the third tier, indicating substantial difficulties in maintaining long and accurate conversations.
@@ 5.2 Dialogue Survival Analysis
To visualize and compare the long-term memory management capabilities of the models over time, we tracked the percentage of active dialogue sessions remaining at each turn, up to a maximum of 50 turns. This yields a dialogue survival curve for each model, as depicted in Figure 3. A performance stratification between the top-3 models and the rest emerges at turn 4, right after the fast exhaustion of the user patience. Initially, MiniMax-M2 demonstrates instruction-following capabilities comparable to GPT-5 and Gemini-3.0-Pro. However, its performance drops dramatically after 10 turns, becoming indistinguishable from second-tier models by turn 15.
@@ 5.3 Fine-Grained Analysis of Constraints
GPT-5 and Gemini-3.0-Pro demonstrate strong, well-rounded performance, topping the rankings on most objective constraints. However, they lag behind Qwen3-235B and Grok-4-Fast on subjective tasks. MiniMax-M2 does not achieve outstanding performance in any single category but ultimately outperforms the remaining models that exhibit spiky profiles.
@@ 7 Limitations
Besides, following previous work, such as Arena-Hard li2024crowdsourced and MT-Bench zheng2023judging, we choose the LLM-as-a-judge approach for subjective constraint evaluation. We acknowledge the inherent limitations of this approach, such as family bias spiliopoulou2025play. In this work, we utilize it as a widely-adopted verifier and prompt it with detailed instructions. Notably, we observed no significant family bias using GPT-4.1, given that it did not disproportionately prefer GPT-5 across subjective tasks. This judge could also be replaced by targeted classifiers. Developing more robust verification methods lies beyond the scope of this paper.
@@ A.1 Basic Metrics
Following prior work zhang-etal-2025-cfbench; li-etal-2025-structflowbench, we quantify overall instruction-following accuracy. Let $N$ be the total number of turns replied by the model. We adopt the following metrics:
@@ B.1 Topic
We collected 541 prompts from IFEval zhou2023instruction. One annotator was tasked with removing the instructions and constraints from each prompt to extract the corresponding dialogue topic. Customized keywords for each topic were then generated by GPT-4.1. Subsequently, each topic and its keywords were verified by two additional annotators. Modifications were iteratively adopted until the data was accepted by both of them.
@@ B.2 Constraint Group
We collected constraints from existing works zhou2023instruction; li-etal-2025-structflowbench and related researches. Utimately, the constraints were classified into 12 groups as shown in Table LABEL:constraint_groups. 9 of them are objective, verifiable with rule-based functions using existing parser packages or regular expressions. The remaining 3 subjective groups draw inspiration from prior work on style transfer heylighen1999formality, emotion recognition busso2008iemocap and age bias analysis liu2024generation. These subjective constraints are measured by adopting GPT-4.1 as a judge, following previous work li2024crowdsourced; zheng2023judging. Specifically, GPT-4.1 is prompted to score constraint satisfaction on a scale of 1 to 10 with detailed explanations. We consider scores greater than 6 as accepted.
@@ B.3 Style
We randomly selected 500 personas from nvidia2025. Then, we employed GPT-4.1 to infer the most plausible language style and tone each persona would use in daily conversation with 3 to 5 descriptive phrases. These phrases serve as inputs to the Query Synthesis Agent to facilitate the generation of diverse and engaging user queries.

### v3

@@ Abstract
Evaluating LLMs’ instruction-following ability in multi-topic dialogues is essential yet challenging. Existing benchmarks are limited to a fixed number of turns, susceptible to saturation and failing to account for users’ interactive experience. In this work, we propose a novel framework featuring a three-layer tracking mechanism and a query synthesis agent to mimic sequential user behaviors. Grounded in Flow Theory, we introduce process-centric metrics and terminate a conversational evaluation only upon exhausting user patience. Leveraging this framework, we present EvolIF, an evolving benchmark covering 12 constraint groups. Our analysis reveals deficiencies in failure recovery and fine-grained instruction following, with performance stratification becoming evident as conversational depth increases. GPT-5 demonstrates the most sustained resilience, maintaining a 66.40% robustness score, outperforming Gemini-3-Pro by 5.59%, while other models lag behind. Data and code will be released at https://github.com/JiaQiSJTU/EvolIF.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has catalyzed the development of increasingly sophisticated applications, ranging from extended conversational systems Rakotonirina et al. (2025) to autonomous agent frameworks Hu et al. (2025). The efficacy of these systems is fundamentally predicated on an LLM’s ability to consistently adhere to instructions throughout conversations spanning multiple topics with evolving constraints. This core capability demands robust long-context processing and stateful memory management. Consequently, designing evaluation frameworks for multi-turn instruction following has emerged as a critical research focus He et al. (2024b); Kwan et al. (2024); Li et al. (2025b).
@@
IFEval Zhou et al. (2023) 1
ComplexBench Wen et al. (2024) 1
Multi-IF He et al. (2024b) 3
@@
Meeseeks Wang et al. (2025a) 3
EIFBENCH Zou et al. (2025) 1
@@
Existing benchmarks suffer from limitations that impede effective evaluation, as exemplified in Fig. 1. First, they fail to capture the interaction dynamics Hao et al. (2024); Zhang et al. (2025a) and extended duration typical of real-wolrd scenarios. As shown in Table 1, most benchmarks are restricted to a short interaction window, predominately fewer than 7 turns Kwan et al. (2024), and neglect scenarios involving interleaved topics He et al. (2024b); Fan et al. (2025). Second, their static nature leads to rapid performance saturation. As LLMs advance, fixed benchmark challenges are quickly mastered He et al. (2024b); Bai et al. (2024). Although some benchmarks offer adjustable complexity Li et al. (2025c), maintaining challenge levels via continuous sample generation incurs prohibitive computational costs for model re-evaluation. Third, current methodologies overlook the process-centric aspects of user experience. Inheriting the paradigm from single-turn tasks Zhou et al. (2023); Zhang et al. (2025b), these benchmarks prioritize final-answer accuracy He et al. (2024b); Li et al. (2025b); Wang et al. (2025a). They neglect interaction stability and fail to provide a direct indication of the maximum number of turns that LLMs can maintain high-fidelity instruction following.
To overcome these shortcomings, we propose a novel and extensible framework for the dynamic generation and process-centric evaluation of complex multi-turn dialogues. Our approach decouples user queries into underlying intentions and surface form. Intention is tracked via a three-layer mechanism that simulates dynamic user behaviors, while the surface form is synthesized by an agent equipped with an LLM-based generator and rigorous validity checkers. Besides, we move beyond single-turn accuracy to emphasize process-centric experience, drawing upon the Flow Theory Csikszentmihalyi and Csikzentmihaly (1990). We introduce the notion of patience to model user stickiness to a conversation, where consecutive frustrations lead to a dialogue termination. And we define a suite of process-centric metrics to quantify user experience such as endurance and robustness.
Leveraging this framework, we introduce EvolIF, a benchmark grounded on 541 topics, 12 groups of commonly-adopted constraint groups and 500 diverse user styles. Through an evaluation of 10 leading LLMs, we observe a distinct performance stratification. GPT-5 and Gemini-3-Pro establish a commanding lead, whose process-centric scores are nearly double or triple of open-source models. Besides, LLMs share a common and steepest performance drop at around turn 5 and 12, revealing critical bottlenecks in their ability to manage accumulated constraints and complex state transitions.
@@ 2.1 Multi-turn Dialogue Benchmarks
First, script-based evaluations Li et al. (2025b); Deshpande et al. (2025); Jia et al. (2025) utilize static conversational logs, derived either from human-bot interactions or simulated histories, to assess a model’s response to the final user query. While this approach ensures controlled and consistent LLM comparison, it fails to capture the interactive nature of dialogue, where a model’s prior responses fundamentally influence the conversational trajectory.
Second, a line of work employs pre-defined templates Zheng et al. (2023); Fan et al. (2025); Han (2025). This approach is labor-intensive, requiring significant human effort to design fixed user query sequences. Consequently, these benchmarks face scalability limitations regarding conversational depth and are susceptible to saturation as models become overly optimized to the test set over time.
Third, researchers have explored using LLMs as user simulators Zhu et al. (2024); Sekulic et al. (2024) and evaluation methods based on conversations between LLMs Duan et al. (2024); Zhao et al. (2025). Nevertheless, such interactions are prone to uncontrolled divergence and exhibit inherent biases, such as family bias Wataoka et al. (2025).
@@ 2.2 Instruction Following Benchmarks
One line of work assesses models’ capabilities within increasingly intricate single-turn interaction. Early benchmarks like CIF Li et al. (2024b) evaluated a single constraint per instruction. Subsequent work has evolved to incorporate multiple constraints Zhou et al. (2023); Jiang et al. (2024); Wen et al. (2024); He et al. (2024a) or multiple tasks Chen et al. (2024); Zou et al. (2025).
A parallel stream of work benchmarks models’ instruction adherence across multiple turns. Multi-IF He et al. (2024b) extends IFEval to 3 turns, while MultiTurnInstruct Han (2025) employs pre-defined templates for diverse scenarios. StructFlowBench Li et al. (2025b) leverages 6 structure types to curate complex dialogue histories. Other studies focus on specialized abilities, such as self-correction Wang et al. (2025a), or domain-specific tasks like code generation Wang et al. (2025b).
@@ 3.1 Overview
Flow theory Csikszentmihalyi and Csikzentmihaly (1990) points out that users enter a psychological state of immersion which can withstand minor disruptions but collapses under prolonged failure. Underpinned by this theory, we employ an adaptive evaluation protocol where the length of a dialogue is contingent on model performance, governed by a “patience” threshold. High-performing models face progressively longer and more challenging threads, while repeated failures will deplete patience and trigger termination. In this way, our benchmark remains a persistent challenge for advanced models, resisting from saturation.
@@ 3.3 Evaluation Protocol
Our evaluation protocol is adaptive and designed to mirror real-world user interactions, premised on Flow Theory Csikszentmihalyi and Csikzentmihaly (1990) and the cooperative principles of dialogue Grice (1975). Repeated failures by a conversation partner serve as a primary catalyst for user frustration, leading to the eventual disengagement Ang et al. (2002); Hernandez Caralt et al. (2025).
@@
$P_{t}=\begin{cases}P_{t-1}-1,&\text{if failed,}\\ P_{max},&\text{otherwise.}\end{cases}$ (3)
@@ 3.4 Evaluation Metrics
Conventional metrics, such as Constraint Satisfaction Rate (CSR) and Instruction Satisfaction Rate (ISR)Zhang et al. (2025b); Li et al. (2025b), focus primarily on outcome accuracy. To capture the nuances of the conversational process, we introduce a suite of process-centric metrics. Given a benchmark of $D$ dialogues, these metrics are defined below (see AppendixA for details).
@@ 4.1 EvolIF Benchmark
We first curated its core assets: topics, constraints and styles. We collected 541 dialogue topics from IFEval Zhou et al. (2023), manually removing the attached constraints to isolate the core task scenarios and subjects. To support our dynamic generation process, we assigned a set of customized keywords for each topic. Concurrently, we consolidated constraints from prior works Zhou et al. (2023); Li et al. (2025b) and our own construction, and systematically re-categorized them into 12 mutually exclusive groups based on semantic intention. These contain 9 objective constraints assessed by rules and 3 subjective constraints measured with an LLM judge. Moreover, we gathered 500 styles by prompting GPT-4.1 with personas from Meyer and Corneil (2025). More in Appendix B.
We guarantee the quality of the benchmark through the following considerations. To ensure integrity and complexity, we applied an automated filter to discard trivial samples, removing dialogues where the average number of constraints over the first 20 turns was less than two. To mitigate family bias Spiliopoulou et al. (2025) introduced by a single synthesizer, we adopted GPT-4.1, Gemini-2.5-Flash and DeepSeek-V3.1 as synthesizers to generate dialogue sessions with $k=3$ trials.
@@ 4.2 Evaluated Models
We conducted evaluation on ten state-of-the-art large language models from different institutions. They include GPT-5-2025-08-07, Gemini-3-Pro-Exp Comanici et al. (2025), DeepSeek-V3.2-Exp Liu et al. (2024a), Kimi-K2-Instruct-0905 Team et al. (2025), Qwen-235B-A22B-Instruct-2507 Yang et al. (2025), Grok-4-Fast-Reasoning, Llama-4-Maverick, Seed-1.6-Thinking-250715, MiniMax-M2 and Mistral-Large-3. All of the models were evaluated with corresponding default settings. Code and data will be released.
@@ 5.1 Main Results
The performance of LLMs on EvolIF is presented in Table 2. Our analysis reveals a distinct stratification in multi-turn instruction-following capabilities. GPT-5 establishes itself as the state-of-the-art, with Gemini-3-Pro following closely behind. These two models demonstrate a superior performance, achieving process-centric scores that are double or even triple of subsequent models. MiniMax-M2 emerges as the most competitive open-source LLMs, forming a second tier alongside Kimi-K2, Qwen3-235B and Grok-4-Fast. The rest constitute the third tier, indicating substantial difficulties in maintaining long and accurate conversations.
@@ 5.2 Dialogue Survival Analysis
To visualize and compare the long-term memory management capabilities of the models over time, we tracked the percentage of active dialogue sessions remaining at each turn, up to a maximum of 50 turns. This yields a dialogue survival curve for each model, as depicted in Figure 3. A performance stratification between the top-3 models and the rest emerges at turn 4, right after the fast exhaustion of the user patience. Initially, MiniMax-M2 demonstrates instruction-following capabilities comparable to GPT-5 and Gemini-3-Pro. However, its performance drops dramatically after 10 turns, becoming indistinguishable from second-tier models by turn 15.
@@ 5.3 Fine-Grained Analysis of Constraints
GPT-5 and Gemini-3-Pro demonstrate strong, well-rounded performance, topping the rankings on most objective constraints. However, they lag behind Qwen3-235B and Grok-4-Fast on subjective tasks. MiniMax-M2 does not achieve outstanding performance in any single category but ultimately outperforms the remaining models that exhibit spiky profiles.
@@ 6 Conclusion
Besides, following previous work, such as Arena-Hard Li et al. (2024a) and MT-Bench Zheng et al. (2023), we choose the LLM-as-a-judge approach for subjective constraint evaluation. We acknowledge the inherent limitations of this approach, such as family bias Spiliopoulou et al. (2025). In this work, we utilize it as a widely-adopted verifier and prompt it with detailed instructions. Notably, we observed no significant family bias using GPT-4.1, given that it did not disproportionately prefer GPT-5 across subjective tasks. This judge could also be replaced by targeted classifiers. Developing more robust verification methods lies beyond the scope of this paper.
@@ A.1 Basic Metrics
Following prior work Zhang et al. (2025b); Li et al. (2025b), we quantify overall instruction-following accuracy. Let $N$ be the total number of turns replied by the model. We adopt the following metrics:
@@ B.1 Topic
We collected 541 prompts from IFEval Zhou et al. (2023). One annotator was tasked with removing the instructions and constraints from each prompt to extract the corresponding dialogue topic. Customized keywords for each topic were then generated by GPT-4.1. Subsequently, each topic and its keywords were verified by two additional annotators. Modifications were iteratively adopted until the data was accepted by both of them.
@@ B.2 Constraint Group
We collected constraints from existing works Zhou et al. (2023); Li et al. (2025b) and related research. Ultimately, the constraints were classified into 12 groups as shown in Table 5. 9 of them are objective, verifiable with rule-based functions using existing parser packages or regular expressions. The remaining 3 subjective groups draw inspiration from prior work on style transfer Heylighen and Dewaele (1999), emotion recognition Busso et al. (2008) and age bias analysis Liu et al. (2024b). These subjective constraints are measured by adopting GPT-4.1 as a judge, following previous work Li et al. (2024a); Zheng et al. (2023). Specifically, GPT-4.1 is prompted to score constraint satisfaction on a scale of 1 to 10 with detailed explanations. We consider scores greater than 6 as accepted.
@@ B.3 Style
We randomly selected 500 personas from Meyer and Corneil (2025). Then, we employed GPT-4.1 to infer the most plausible language style and tone each persona would use in daily conversation with 3 to 5 descriptive phrases. These phrases serve as inputs to the Query Synthesis Agent to facilitate the generation of diverse and engaging user queries.